OpenStudy (anonymous):

Find the Taylor series for \[f(x)= \frac 1x; a=-3\] \[f(x)=\sum_{n=0}^{\infty} \frac{f^n (-3)}{n!} x^n= \frac{1}{-3}-\frac{\frac 19}{1!}x-\frac{\frac{1}{27}}{2!}x^2-\frac{\frac{1}{81}}{3!}x^-\frac{\frac{1}{243}}{4!}x^4-.....\] how did I do?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

Whaaaaaatt?

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

<tilting head>

OpenStudy (anonymous):

\[ f(x)=\sum_{n=0}^{\infty} \frac{f^n (-3)}{n!} (x+3)^n= \frac{1}{-3}-\frac{\frac 19}{1!}(x+3)-\frac{\frac{1}{27}}{2!}(x+3)^2\cdots \]

OpenStudy (anonymous):

where did you find the (x-3)?

OpenStudy (anonymous):

This the definition of the Taylor Series as a =-3 \[ f(x) =\sum_{n=0}^\infty\frac {f^{n}(a)}{n!}(x-a)^n \]

OpenStudy (anonymous):

was I doing the MacLaurin series at (-3) ?

OpenStudy (anonymous):

sorry I'm new to Maclaurin, Taylor, and Co.

OpenStudy (anonymous):

Ok so if it is not centered at around zero I use the Taylor series?

OpenStudy (anonymous):

Well the question asked for Taylor series...I'm soo silly

OpenStudy (anonymous):

\[ -\frac{1}{3}-\frac{x+3}{9}-\frac{1}{27} (x+3)^2-\frac{1}{81} (x+3)^3-\frac{1}{243} (x+3)^4-\frac{1}{729} (x+3)^5-\frac{(x+3)^6}{2187}-\\\frac{(x+3)^7 }{6561}-\frac{(x+3)^8}{19683}-\frac{(x+3)^ 9}{59049}-\frac{(x+3)^{10}}{177147}+O\left ((x+3)^{11}\right) \]

OpenStudy (anonymous):

the denominator goes to infinity

OpenStudy (anonymous):

If you notice that \[ f^n(x)=(-1)^n \frac { n!}{x^{n+1}} \]

OpenStudy (anonymous):

then \[ \frac {f^n(a)}{n!}(x-a)^n =(-1)^n \frac {(x-a)^n}{a^{n+1}} \] Then the series will be \[ \sum_{n=0}^\infty (-1)^n \frac {(x-a)^n}{a^{n+1}} \]

OpenStudy (anonymous):

give me a min to take that in...

OpenStudy (anonymous):

for a =-3 \[ \sum_{n=0}^\infty (-1)^n \frac {(x+3)^n}{(-3)^{n+1}}=\sum_{n=0}^\infty - \frac {(x+3)^n}{3^{n+1}} \]

OpenStudy (anonymous):

makes sense. Remind me what the first equation is \[f^n(x)=(-1)^n \frac { n!}{x^{n+1}}\]

OpenStudy (anonymous):

Power series?

OpenStudy (anonymous):

I remember doing a ratio test with \[f^n(x)=(-1)^n \frac { n!}{x^{n+1}}\] to test for convergence....alternating series?